According to an​ airline, flights on a certain route are on time 80​% of the time. Suppose 24 flights are randomly selected and the number of​ on-time flights is recorded. ​(a) Explain why this is a binomial experiment. ​(b) Find and interpret the probability that exactly 16 flights are on time. ​(c) Find and interpret the probability that fewer than 16 flights are on time. ​(d) Find and interpret the probability that at least 16 flights are on time. ​(e) Find and interpret the probability that between 14 and 16 ​flights, inclusive, are on time.
Answer #1

Solution-:

Given: n=24,p=0.80

Let, X- be the number of on-time flights is recored among the 24 flights

(a) This situation is siutable for binomial distribution with parameters n=24 and p=0.80

\therefore X\rightarrow B(n=24,p=0.80)  

The p.m.f. of X is given by,

24 P(X = z) = *p*(1 - p)24-- CI= 0, 1, 2, .24

:0 <p< 1;q = 1 — р

  =0;Otherwise

We find all probability in following table:

n=24 p=0.80 q=0.20
x nCx p^x q^(24-x) P[X=x]
0 1 1 1.67772E-17 0.0000
1 24 0.8 8.38861E-17 0.0000
2 276 0.64 4.1943E-16 0.0000
3 2024 0.512 2.09715E-15 0.0000
4 10626 0.4096 1.04858E-14 0.0000
5 42504 0.32768 5.24288E-14 0.0000
6 134596 0.262144 2.62144E-13 0.0000
7 346104 0.209715 1.31072E-12 0.0000
8 735471 0.167772 6.5536E-12 0.0000
9 1307504 0.134218 3.2768E-11 0.0000
10 1961256 0.107374 1.6384E-10 0.0000
11 2496144 0.085899 8.192E-10 0.0002
12 2704156 0.068719 4.096E-09 0.0008
13 2496144 0.054976 2.048E-08 0.0028
14 1961256 0.04398 1.024E-07 0.0088
15 1307504 0.035184 0.000000512 0.0236
16 735471 0.028147 0.00000256 0.0530
17 346104 0.022518 0.0000128 0.0998
18 134596 0.018014 0.000064 0.1552
19 42504 0.014412 0.00032 0.1960
20 10626 0.011529 0.0016 0.1960
21 2024 0.009223 0.008 0.1493
22 276 0.007379 0.04 0.0815
23 24 0.005903 0.2 0.0283
24 1 0.004722 1 0.0047

(b) We find, P[exactly 16 flights are on time]

=P[X=16]

=0.0530 (Using above table)

Therefore, the required probability is 0.0530.

This is interpreted as about a 5.30\% chance is exactly 16 flights out of 24 being on time.

(c) P[that fewer than 16 flights are on time]

  =P[X<16]

=P[X\leq 15]

=P[X=0]+P[X=1]+P[X=2]+P[X=3]+P[X=4]+P[X=5]+P[X=6]+P[X=7]+P[X=8]+P[X=9]+P[X=10]+P[X=11]+P[X=12]+P[X=13]+P[X=14]+P[X=15]

  =0+0+0+0+0+0+0+0+0+0+0+0.0002+0.0008+0.0028+0.0088+0.0236

(Using above table)

=0.0362

Therefore, the required probability is 0.0362.

This is interpreted as about a 3.62\% chance of fewer than 16 flights out of 24 being on time.

(d) P[at least 16 flights are on time]

=P[X\geq 16]

  =1-P[X< 16]=1-\{P[X=0]+P[X=1]+P[X=2]+P[X=3]+P[X=4]+P[X=5]+P[X=6]+P[X=7]+P[X=8]+P[X=9]+P[X=10]+P[X=11]+P[X=12]+P[X=13]+P[X=14]+P[X=15]\}

=1-0.0362 (using (b) )

  =0.9638

Therefore, the required probability is 0.9638.

This is interpreted as about a 96.38\% chance at least 16 flights out of 24 being on time.

(e) P[14 and 16 ​flights are on time]

  =P[14\leq X\leq 16]

  =P[X=14]+P[X=15]+P[X=16]

=0.0088+0.0236+0.0530

  =0.0854

This can be interpreted as about a 8.54\%\approx 9\% % chance that between 14 and 16 (inclusive of both 14 and 16) flights are on time.